Assume a consumer has $40 to spend and for both products the marginal utilities are shown in the following table: Quantity MUA MUB 1 35 80 2 20 40 3 12 18 Assume that each product sells for $10 per unit. a) How many units of each product will the consumer purchase? b) Assume the price of product B rises to $20 per unit. How will this consumer allocate her budget now? c) If the prices of both products rise to $20 per unit, what will be the budget allocation


1
Expert's answer
2021-11-08T16:43:33-0500

Solution:

A.). Utility quantity maximizing is where: MUA/PA = MUB/PB

Budget constraint: I = PAA + PBB

40 = 10A + 10B

40 = (10 "\times" 2) + (10 "\times" 2)

40 = 20 + 20

40 = 40

To maximize utility, the consumer will purchase two units of product A and two units of product B.

Product A = 2 units

Product B = 2 units


B.). The new budget constraint: 40 = 10A + 20B

To maximize utility, the consumer will now purchase two units of product A and 1 unit of product B.

Product A = 2 units

Product B = 1 unit

40 = 10A + 20B

40 = (10 "\times" 2) + (20 "\times" 1)

40 = 20 + 20

40 = 40


C.). New budget constraint = 40 = 20A + 20B

To maximize utility, the consumer will now purchase 1 unit of product A and 1 unit of product B.

Product A = 1 unit

Product B = 1 unit

40 = 20A + 20B

40 = (20 "\times" 1) + (20 "\times" 1)

40 = 20 + 20

40 = 40